### v1

# ScrollTimes: Tracing the Provenance of Paintings as a Window into History

Abstract
Digital humanities research has flourished due to the diverse artifacts available in cultural heritage databases. However, over-reliance on a single artifact type can result in poor contextualization and a constrained understanding of historical context. We collaborated with art historians to examine handscrolls, a form of traditional Chinese painting which offers a wealth of data for historical analysis and provides a unique opportunity for understanding history through artwork. We propose ScrollTimes, a visual analysis system for tracing handscroll historic context by linking multiple data sources. Specifically, a unique layout is developed for efficiently viewing long handscrolls. Using image processing techniques and language models, we extract, verify, and supplement elements in handscrolls with different cultural heritage databases. Furthermore, interactive biographies are constructed for handscrolls to uncover their historical narratives, provenance trajectories, and artistic legacies. Validated through case studies and expert interviews, our approach offers a window into history, fostering a holistic understanding of handscroll provenance and historical significance.
keywords:
traditional Chinese painting, handscroll, painting biography, neural networks, seam carving
[…]
The visual analysis system ScrollTimes enables art historians to analyze handscrolls by combining multiple databases and constructing biographies of paintings that fit their research interests. In the case of exploring the renowned handscroll “Autumn Colors on the Que and Hua Mountains,” historians first utilized the Handscroll Features View (A) to understand the handscroll elements, such as the core painting, seals, and inscriptions. Then, they further explored the elements of interest in the Seals and Inscriptions View (B) and validated the information using the Linked Data View (B4). Finally, they constructed a biography of the painting that matched their research interests using the Biography View (C).
1 Introduction
The emergence of cultural heritage databases has significantly propelled the development of Digital Humanities (DH). This advancement has transformed traditional research workflows by streamlining manual data collection and annotation, thus enabling scholars to analyze vast amounts of data more effectively [65, 18]. Various DH studies have employed cultural heritage databases that collect only a single artifact type (e.g., ancient literature [69, 68], paintings [39], and sculptures [66]). Although different types of cultural artifacts inherently originate from specific historical contexts and often recount the same historical events, they exhibit diverse themes, formats, meanings, and purposes. Relying on only one artifact type results in a restricted understanding of history [27]. For example, computational painting analysis tends to focus more on the aesthetic value of the painting itself (e.g., research on the style [39], patterns [26], and creation processes [59]), while largely neglecting the painting’s historical value as an artifact.
In this study, we explore the “handscroll” [45], a distinguished format in traditional Chinese painting, as the primary artifact type to reveal the historical context embedded in these artworks, together with multiple data sources. The handscroll is selected because of its unique creation process, which serves as a rich source of historical information over a long period of time. The core painting (Fig. 1B, usually landscapes or narrative stories) is typically drawn on a long piece of silk or paper from right to left. This wide aspect ratio enables the painter to depict the flow of time and the variations in space upon a single handscroll. Upon the completion of the core painting, the painter often adds inscriptions (Fig. 1A, text describing the motivation behind the creation) and seals (Fig. 1C, vermillion imprint indicating their signature). At this point, a handscroll is “born”. However, its development does not cease, as the “life” of the handscroll continues. As future generations appreciate the handscroll, they frequently contribute additional content, such as mounting a new frontispiece or colophon with their own inscriptions (including critiques or historical event records) and seals (indicating appreciation). Therefore, the handscroll not only records history but also becomes a part of it [28].
Figure 1: Basic handscroll format. It is a long piece of paper or silk that can be rolled up or displayed, and viewers can see different parts of the painting when it is opened from right to left.
The centuries-old handscrolls resonate with historic significance, and art historians have been studying their provenance over time based on their rich components. We propose creating biographies for handscrolls to explore their historical narratives. By recording their “life stories” (i.e., appreciation records and anecdotes) and exploring the events involving figures who interacted with the paintings, we aim to use the paintings to verify historical events, providing new perspectives and insights into the longitudinal study. In addition, since traditional Chinese painting places great emphasis on copying and transferring styles [16], the inheritance of painting styles can be traced through comparisons of similarities, akin to viewing the “genealogy” of a painting (i.e., other paintings that were influenced by it). The relationships between painters can also be discovered by linking with other cultural heritage databases, thereby offering cross-sectional data to broaden the scope of historical research.
In collaboration with art historians, we have identified three primary challenges in building biographies of handscrolls: 1) the difficulty in viewing handscrolls due to their wide aspect ratios, 2) the inefficiency of conventional methods in extracting and analyzing inscriptions and seals to understand the provenance of paintings, and 3) the overwhelming task of aggregating multiple data sources, which relies on serendipity [65] if prior knowledge is not incorporated into the workflow.
To enhance the viewing experience, we leverage seam carving [9] to reduce the white space in handscrolls. Then the wide and narrow paintings are transformed into a circular format, enabling more efficient, interactive viewing for art historians. To capture essential components, we employ deep neural networks [14, 31, 32] for extracting core painting and seals, and then retrieving implicit information from multiple databases. Moreover, we utilize an NLP model [63] to extract entities such as time, location, figures, and things from inscriptions in Classical Chinese, then multiple databases are used to verify existing data and supplement missing data. Finally, we propose the ScrollTimes, an interactive visual analysis system to integrate multiple databases and offer validation and supplementation from various perspectives, assisting art historians in comprehending painting biographies and analyzing the provenance history of the paintings. In summary, this study has the following contributions:
  • We developed a new painting viewing method, the circular view, based on handscroll characteristics, enabling art historians to swiftly view paintings.
  • We used automatic methods and multiple databases to extract, validate, and supplement elements in paintings, allowing art historians to create painting biographies interactively using ScrollTimes.
  • We conducted case studies and expert interviews to demonstrate the effectiveness and usability of our approach.
2 Related Works
Our research is closely aligned with the topics about the visualization of linked data for cultural heritage and computational methods for analyzing paintings. Consequently, we will now delve into the body of literature surrounding these fields.
2.1 Visualize Linked Data for Cultural Heritage
With the advent of digital cultural heritage databases around the world, open linked data provides a formidable foundation for reconstructing the historical evolution of culture [64, 65, 29, 52, 21]. Researchers leverage various visual analysis techniques such as Semantic Web technology [19, 43, 24, 29], multi-temporal and spatial reconstruction [53, 12], and web-GIS [50] to structure and elegantly display the heterogeneous data. These works largely promote resource sharing and comprehensive information exploration.
Researchers are increasingly cognizant of the significance of cultural heritage databases [24, 25, 65, 54]. For instance, some studies use semantic web technology and natural language processing methods to enhance access to cultural information and address the public challenges of managing small collections. Meanwhile, visualization methods for digital cultural heritage collections have been scrutinized and reviewed, and new applications (such as “Casual Learn” [57]) and tools (such as CHEST [30] and “InTaVia” [48]) have been introduced. The “InTaVia” project has developed a comprehensive visualization approach for a multi-modal historical knowledge graph while emphasizing the importance of a user-centered design process. Nevertheless, the credibility of the visual analysis results for cultural heritage remains a challenge due to the intricacy and diversity of cultural heritage data, which often engenders ambiguity and misunderstanding. Additionally, current platforms lack better integration of prior domain expert knowledge, which plays a crucial role in professional usages.
Our system endeavors to furnish art historians with validated entity information from multiple data sources, integrates the information for art historians to make inferences with prior knowledge, and ultimately offers easy-to-use interactions to refine the information based on requirements.
2.2 Computational Approaches for Analyzing Painting
Recently, a surge of interest has emerged in constructing painting databases from various angles, including the painter [42], theme [46, 37], and style [40] of paintings. Some of them support cross-modal retrieval by combining text and image features [23, 47]. These databases have greatly supported the interpretation [27, 26], creation [61, 67], and exhibition [36, 41] of traditional Chinese painting. However, these works tend to focus more on the aesthetic value of the painting itself. MTFFNet [39] has been proposed to analyze the painting style of painters, Fanet al. [26] focus on discovering painting patterns, and Tanget al. [59] reproduce the painting process for better painting education. The historical value of paintings remains unexplored.
In fact, behind these paintings is huge historical information on politics [13], economy [34], and culture [17]. Several studies have presented fundamental historical background information of paintings through diverse methods. For instance, “Hieronymus Bosch,” [3] presents art history through online multimedia, and “Mapping Paintings,” [4] employs maps to display provenance records. However, they are limited to neither engaging users’ prior knowledge nor providing comprehensive information. Besides, David et al.’s [20] study on the attribution of paintings also lacks sufficient data verification techniques. These works may not fully meet the research needs of art historians who seek to explore the deeper historical context of paintings.
This study aims to explore the historical value of Chinese handscroll paintings and develop a visual analysis system tailored to the needs of art historians. To this end, we employ automated methods to extract handscroll elements and validate them by cross-referencing with relevant historical contexts from external cultural heritage databases. We then construct biographies for the paintings and utilize event sequence visualization techniques [33, 44] to comprehensively demonstrate the historical significance.
3 Background
This section provides an overview of the domain background and outlines the requirements obtained through in-depth interviews with domain experts. Over the past year, we have collaborated closely with five highly experienced art historians. This group includes three professors (A1-A3), each possessing over twenty years of research experience in the field of traditional Chinese painting history, and two Ph.D. students (A4 and A5) with nine years of practical experience in traditional Chinese painting. In addition, we seek guidance on cultural heritage data processing from two editors (E1 and E2) of the China Biographical DataBase [11]. Our collaboration took place in three stages: task analysis (Section 3.2), system design (Section 5), and system evaluation (Section 6). We first present the data description in Section 3.1.
3.1 Data Description
This study collects 1380 handscroll data (HD) from the open dataset platform [49]. Each handscroll contains the following data elements:
  • Core painting (Fig. 1B): main content of the handscroll, often depicting a narrative.
  • Inscriptions (Fig. 1A): texts on the handscroll, containing painters’ records, historical events, and critiques of the painting.
  • Seals (Fig. 1C): red seal stamps imprinted on the handscroll by the owner or appreciator, indicating signature or authentication.
However, these data elements are usually ambiguous and incomplete, which hinders art historians’ systematic analysis. To supplement the handscroll data and perform cross-validation, we selectively collect data from five databases containing information on Chinese historical figures, time periods, locations, and events. These databases (Fig. 2) are chosen based on their relevance to the study of Chinese painting history and their reputation within the academic community. The selected databases include CBDB, CHFSD (Chinese Historical Figure Seal Data) [2], PerAD (Person Authority Database) [5], PlaAD (Place Authority Database) [6], and TAD (Time Authority Database) [7].
Figure 2: The Handscroll Database is validated and supplemented through the integration of data from multiple data sources.
3.2 Requirement and Task Analysis
We interviewed our collaborators to gain insights into their research process and inform our system design. Accordingly, we identified the system requirements and design tasks necessary to address the challenges that art historians have been facing when performing conventional analysis using current techniques.
R1: Efficiently browse handscroll and capture its features. Handscrolls’ wide aspect ratio makes traditional study methods challenging, as art historians need to repeatedly scan the handscroll to connect the core painting, inscriptions, and seals to gain insights. A5 indicates that this may take up to 10 hours. Our system should provide an efficient solution for quickly browsing and capturing handscroll features.
  1. T1
    Provide interactive browsing methods to optimize the viewing of handscrolls. Digitization of cultural artifacts offers coarse-to-fine viewing of painting details. However, handscrolls’ wide aspect ratios and excessive white spaces pose challenges to traditional viewing methods. The handscroll viewing experience shall be optimized with convenient and interactive browsing methods.
  2. T2
    Extract and visualize the features of handscrolls. Handscrolls contain elements with rich information, and there exist latent correlations among these elements. However, manual acquisition and connection are time-consuming. To improve the efficiency of handscroll analysis, we intend to utilize computer technology to extract features and visualize the correlations.
R2: Validate and supplement handscroll features from multiple perspectives automatically. To enhance the credibility of analysis results, art historians habitually cross-reference a handscroll’s elements from multiple perspectives to reduce ambiguity and incompleteness. And additional historical context is required to attain a comprehensive understanding of the handscroll. Our system should offer an efficient and effective approach to releasing historians from heavy labor.
  1. T3
    Cross-reference handscroll features from multiple databases. A1 points out that the traditional analysis workflow often necessitates perusing extensive literature across multiple platforms to cross-reference handscroll elements, which is time-consuming. It is imperative to establish automated methods that employ various databases for cross-references handscroll elements, improving the reliability of art historians’ analysis.
  2. T4
    Provide rich historical contextual information for handscroll features. Art historians not only need to connect the elements within the artwork but also relate to external information, which can provide them with new research perspectives and insights. However, relying on a single data source cannot achieve this. We need to combine different cultural heritage databases to provide rich historical contextual information.
R3: Construct painting biographies according to user’s requirements interactively. Art historians have varying standards for constructing biographies of paintings based on their own research interests. For example, A4 is skilled at creating rigorous and completed painting biographies primarily from a literary perspective, while A5 hopes to freely explore the biographies based on the similarity of core paintings. After automatically generating the painting biographies, our system should further enable art historians to customize the painting biographies from different perspectives in an interactive manner.
  1. T5
    Refine painting biographies interactively. Historical data in databases is often limited by data sources, e.g., rare literary documents, which can result in missing or incomplete information in painting biographies. Interactive supplementation by art historians can effectively address this issue. The system provides data from multiple sources, and art historians can use their prior knowledge to integrate and supplement the data elements. This interaction can help to create painting biographies that are comprehensive and more aligned with their interests.
  2. T6
    Adjust painting biographies accordingly. A1 and A3 indicate that art historians have different interests in the information level of painting biographies. Typically, the content within the painting itself, the complementary information from external data sources, and the cross-data recommendations from automatic computer technologies are the three levels that art historians are mostly concerned with. Therefore, art historians should be enabled to interactively adjust different information levels of the painting biography accordingly.
4 Model
This section introduces the techniques utilized for information extraction and processing of various elements in the handscroll, including inscriptions, core painting, and seals. Furthermore, it elaborates on the rigorous process of verifying and supplementing the extracted information with multiple databases to obtain a comprehensive understanding of the events depicted in the handscroll.
4.1 Handscroll Feature Extraction
We commence our approach by leveraging image processing and natural language processing techniques to extract pertinent information from the handscroll (T1).
4.1.1 Inscription Processing
The inscriptions used in this work come from the HD database. We directly employ the Named Entity Recognition (NER) technique to process inscriptions’ textual contents (Fig. 3A). Specifically, NER aims to identify entities with specific meanings in the text, including person names, place names, organization names, and proper nouns. We adopt a Chinese NER model called RaNER [63], which incorporates a Transformer-CRF model with StructBERT [62] as the pre-trained model. This model can identify entities such as addresses, people, times, organizations, and other proper nouns in the inscriptions.
Nevertheless, NER models are subject to limitations regarding the length of the input text, hindering the processing of some inscription texts. To alleviate the obstacle, we deliver a sliding window approach to extract sentences of reasonable length, which are then fed into the model for entity recognition. Subsequently, a straightforward voting strategy is applied to manage the overlapping short sentence parts, culminating in the final NER classification results. To obtain additional details regarding entities like figures, times, and locations, the NER results are post-processed to associate them with open-source databases. However, the model tends to recognize boundaries that are excessively large, leading to erroneous results during the testing phase. To address this issue, a sliding window is implemented that is based on individual characters to extract sub-words, which are subsequently queried separately. If a sub-word produces a relevant result, the range of the recognition results is rectified and associated with the corresponding information. This approach effectively boosts the accuracy of entity recognition, as demonstrated by test results.
Figure 3: After being processed by various models and validated and supplemented through multiple databases, inscription (A), core painting (B), and seal (C) are organized into a series of historical events (D) for the purpose of constructing the painting biography through the ScrollTimes (E).
4.1.2 Core Painting Processing
The relationship between students and teachers among painters plays a crucial role in shaping their painting techniques and styles. To provide an insightful exploration of a large number of paintings, it is necessary to recommend similar paintings. As a handscroll usually contains a substantial amount of textual content, it is essential to extract the core painting from the handscroll before calculating similarities, ensuring a more precise and rigorous calculation of painting similarity (Fig. 3B).
In this paper, we utilize a deep neural network known as DeepLab V3 [14] to extract core painting. DeepLab V3 has achieved impressive performance in semantic segmentation within the field of natural images. It uses the ResNet [35] as the backbone for feature extraction, complemented by an Atrous Spatial Pyramid Pooling module. This module analyzes high-level features extracted from the image and generates multi-scale context features, thereby enhancing the ability to segment objects accurately. After training DeepLab V3, we employ it to extract core paintings in HD, building a core painting gallery.
To measure the similarity between core paintings, we use a self-supervised learning method called BYOL [32]. This method involves projecting images into a feature space using an encoder and computing their similarity in that space. During training, two augmented views of each image are fed into an online encoder and a momentum encoder to construct consistency constraints and optimize the encoder. We then retrieve the most similar core paintings from the gallery by comparing the cosine similarities of their query painting features.
4.1.3 Seal Processing
To detect and match the seals in the handscroll, and understand the owner and content of the seals, we employ computer vision methods for relevant processing (Fig. 3C). We employ a neural network named YOLOX [31] to detect seals in the handscroll. YOLOX integrates the benefits of a series of YOLO models, offering exceptional detection speed and accuracy. In the experiments, we curate a dataset of Dynasty Song paintings and perform annotations on the seals within the paintings. The trained model facilitates the detection of seals from handscrolls and their subsequent isolation for seal matching. The handscrolls contain numerous seals, and their attribution to specific authors or content may prove challenging. Thus, matching with available open-source seal databases becomes necessary. We collect a standard seal gallery including 36302 seal images, along with the corresponding author and textual content. By leveraging this gallery, we match the detected seals from handscrolls to issue the author and content information to the seals.
We train another BYOL model with a ResNet-101 backbone for seal feature extraction. Notice that seal matching performance may be impacted by the intricate and diverse backgrounds of the handscrolls. Moreover, the gallery samples are limited, with each seal in the gallery having only one image. Therefore, we augment the gallery by gathering a set of seal backgrounds from handscrolls and randomly substituting the seal backgrounds to create a new seal dataset with several times more samples. To replace seal backgrounds, we apply Otsu’s method [51] for binary thresholding, which is efficient and effective. With the trained BYOL, we project seal images into feature vectors for seal matching. For each detected seal from handscrolls, we employ a vector retrieval algorithm, specifically Local Sensitive Hashing [38], to search for the most similar seals from the gallery, thereby accelerating the retrieval process.
4.2 Data Fusion
In Section 4.1, we extracted entities, including time, location, figure, and thing from inscriptions core painting, and seals, but the accuracy and consistency of the information have not been validated. To ensure a clear and well-structured account of events, we cross-check and enrich the data by referring to multiple databases (Fig. 3D) (T3, T4).
Figure validation supplement. To identify the figure entities, we query PerAD and CBDB to validate and supplement the information. However, the identification of the individuals can be challenging due to the extensive use of aliases and personal names with prefixes in ancient Chinese. To address this issue, we adopt the following approaches:
  • Directly querying PerAD and CBDB. Although CBDB has comprehensive descriptions of historical figures, PerAD provides valuable data, especially related to religious figures and deities. For example, the painter Zhao Mengfu can be queried in both databases, but the named entity Jin Mu can only be found in PerAD as an alternative name for Queen Mother of the West [10].
  • Querying with sliding-window segments. For instance, “Han Lin Qian Pu”(“Hanlin” was an academic and administrative institution of higher learning founded in China) can be segmented and corrected as Qian Pu to supplement the information of the corresponding figure.
  • Priority determination for figures with the same name based on the era and social relationships. For example, an inscription from the Yuan Dynasty mentions a figure named Gong Jin Fu, where “Fu” is an honorific for a man. Despite there being more than twenty historical figures with the name Gong Jin, only Zhou Mi and Li Kezhong is from the Yuan Dynasty, with 47 and 3 social relationships, respectively. Therefore, we can conclude that the figure referred to is Zhou Mi.
  • Allowing art historians to interactively choose the correct figure. While automated methods are helpful, they may not guarantee accuracy in identifying historical figures. Therefore, art historians can conduct fuzzy searches on figure names and select the correct figure based on the presented list and their prior knowledge.
Location validation. Place entities have fewer errors than figure names, yet challenges still emerge due to the prevalence of aliases, prefixes, and suffixes. Analogous to the procedure for figures, we cross-reference PlaAD and CBDB databases to rectify place entities. For instance, the toponym Qizhou can be recognized by both databases, while Que mountain can only be found in PlaAD.
Time standardization. The model’s recognition of time entities is generally precise, with the primary task of converting them into Gregorian calendar years. In ancient China, years are typically indicated by the era name followed by either the year number under that era or the Sexagenary cycle [55], a traditional time reckoning system. Regular expression matching is used to extract the corresponding era name, numeric value, and Sexagenary cycle name, and the era’s start and end dates are queried to calculate the exact year. For example, if the start year of the “Yuanzhen” era is 1295, then the time entity Yuanzhen second year corresponds to the year 1296. Qianlong Wuchen denotes the year 1748, as the “Qianlong” era ranges from 1736 to 1796, with 1748 being the “Wuchen” (a Sexagenary cycle name) year.
Thing supplement. Due to the diverse entity types, such as mountains, birds, and paintings, automatic methods face difficulties verifying and supplementing them. Art historians’ prior knowledge is required to supplement events with entities. For example, if a CBDB records that someone wrote inscriptions in paintings, and the art historian adds the entity “Autumn Colors on the Que and Hua Mountains” to the event, it constitutes a complete event: someone inscribes an inscription on“Autumn Colors on the Que and Hua Mountains” painting. As a quantitative analysis, we examine the performance of NER on a separate test set. The post-processed NER results achieve an F1 score of 0.558, significantly surpassing the F1 score of 0.483 yielded by the vanilla NER results.
5 Visual Analytic System
Figure 4: The original painting (A) is transformed into a circular ring shape (B). The seam carving algorithm is used to remove white space (C). After preserving the core painting, reducing silk, and repairing text distortion, the final representation is displayed (D). * To better visualize the silk texture, we have applied visual enhancement to the silk in this caption.
We propose the visual analysis system ScrollTimes to display the features of the handscroll and support art historians in exploring interested figures interactively, helping art historians construct painting biography.
5.1 Handscroll Features View
The Handscroll Features View offers art historians an accessible way to browse handscrolls and gain insights from feature visualization (Fig. 3E1) (T1, T2). The redesign of the handscroll considered two principles: 1) conserving screen space to emphasize the crucial points of the scroll (Section 5.1.1); 2) conforming to contemporary reading habits by modernizing the ancient text reading form (Section 5.1.2).
5.1.1 Handscroll Circle Design
Transforming the handscroll into a circle. To maximize the efficient use of system space for handscrolls with wide aspect ratio disparities (Fig. 4A), we transform them into a circular ring shape (Fig. 4B).
The handscroll is divided into two equal-width parts and projected into a 180-degree semicircular ring, which is then flipped and concatenated to form a complete circular ring. To achieve the desired circular effect, each part of the original image is divided into rectangular blocks and proportionally projected into trapezoids based on pixel positions.
Processing white space. We use seam carving [9, 22] algorithm for handscroll white space processing (Fig. 4C). This involves scaling the original handscroll while preserving high-energy content and removing low-efficiency information. Seam carving constructs an energy map of the painting by calculating the discrete pixel gradients. A seam is defined as a connected path of pixels running from the top to the bottom of the painting. Using dynamic programming, the algorithm finds and removes the connected path with the lowest cumulative energy.
Regulating the ratio of the core painting. In the handscroll, the core painting contains more content and has higher information density than other parts. During interviews with art historians, it is determined that the core painting should occupy one-third of the entire length of the handscroll. We resize the handscroll to a fixed length, with one-third being the extracted core painting (Fig. 4D) and the remaining filled with other parts. In addition, some texts are severely deformed due to excessive compression. To solve this problem, we divide non-core painting parts into smaller blocks and reduce the compression ratio for blocks with over-deformable texts (Fig. 4D).
Reducing silk space. Silk is an inherent part of the handscroll (Fig. 1), but contains relatively less information. However, the complex texture and pattern in the silk material may yield high energy values when calculating gradients, which results in poor squeezing effects on the silk region. Therefore, we introduce the Frequency-tuned (FT) salient region detection algorithm [8], which efficiently detects salient objects with clear boundaries and captures the key boundary information of the painting to preserve important content. By linearly combining the FT saliency map and the gradient map, the energy map calculation process is optimized (Fig. 4D). Specifically, the FT algorithm calculates the saliency map based on the Euclidean distance between the pixel colors and the average color in the Lab color space. The input image $I$ is first Gaussian-filtered to remove high-frequency information, resulting in $I^{\prime}$. With the mean Lab color $\bm{\mu}$ of the original image, we obtain the saliency value in the position $(x,y)$ by $S_{FT}(x,y)=\|\bm{\mu}-\bm{c}(x,y)\|_{2}^{2},$ where $\bm{c}(x,y)$ is the Lab color of the filtered image $I^{\prime}$. We normalize the saliency maps $\bm{S}_{FT}$ and $\bm{S}_{grad}$ obtained from the FT algorithm and the gradient operator, respectively, and then linearly fuse them to generate a normalized energy map as $\bm{S}=\alpha N(\bm{S}_{grad})+\beta N(\bm{S}_{FT}),$ where $N(\cdot)$ is a $\max-\min$ normalization function, and $\alpha$ and $\beta$ are weighting coefficients. We empirically set $\alpha=0.7$ and $\beta=0.3$.
Interaction. We provide three interactive ways for art historians to observe the parts of their interests in (Fig. 4D). 1) [Uncaptioned image] When the mouse moves over the handscroll, the original image of the current section is displayed. 2) [Uncaptioned image] When the mouse slides over the blank space of the circular ring, the ring can be rotated accordingly. 3) [Uncaptioned image] The original handscroll is provided below the circular ring, which supports the art historians in selecting a part of the area. The perspective center of the circular ring will rotate to the selected area. Art historians also can zoom in on the original painting to view the details of the handscroll [Uncaptioned image] .
5.1.2 Handscroll Features Word Cloud
Seals and inscriptions on handscrolls are not systematically distributed and often pose reading challenges for art historians due to the ancient Chinese writing style. To make them more accessible, we apply a word cloud visualization technique to convert ancient writing into a modern format, which facilitates easy access to these features while aligning with modern reading habits. In Section 4.1.3, we extracted the sealers on the handscroll and simplified their visual form into squares (Fig. 3E1-1). The text in the middle of the square represents the sealer’s surname, while the transparency of the red background of the text encodes the number of seals they left on this scroll. This allows art historians to quickly understand which sealer left the most seals on the painting scroll. The color shade of the outer frame reflects the number of seals they left on the entire handscroll dataset, providing insight into their importance in the history of collecting and appreciation (Fig. 5). To reflect the changes in dynasty during the transmission of the painting, we cluster the sealers according to their respective dynasties (Fig. 3E1-1). In Section 4.1.1, we conducted entity extraction on the inscriptions and generated word cloud visualizations based on entity classification ([Uncaptioned image]Time, [Uncaptioned image]Location, [Uncaptioned image]Figure, [Uncaptioned image]Thing). We use font color depth to encode the frequency of the words (Fig. 3E1-2). Upon clicking on a particular sealer, the solid lines on the view illustrate the distribution of their seals on the handscroll, while the dotted lines indicate the entities mentioned in their inscriptions. Additionally, the Seals and Inscriptions View (Fig. 3E1-3) provides a comparison between the seals on the handscroll and the images in the seal database, along with the original inscriptions that have been identified and labeled by the computer based on entity classification. This view is designed to assist art historians in identifying and revising information based on prior knowledge (T3).
5.2 Linked Data View
To provide supplementary information for the handscroll and enable art historians to have a more comprehensive understanding of the painting and conduct relational analysis, we visualize the linked data from other databases (Fig. 3E2) (T4). We select the ego network relationship diagram, which is familiar to art historians, to visualize figure social relationships (Fig. 3E2-1). Additionally, we also present the relationships between cohorts through matrix visualization (Fig. 3E2-2).
5.2.1 Ego Network View
The Ego Network View provides a social network analysis that focuses on a single historical figure. If an art historian is interested in a particular figure from a seal or inscription (Fig. 3E2-1), they can obtain additional information through this view. We encode the data related to the paintings using a hairbrush icon ([Uncaptioned image]Painter) to represent the painter, with the size of the icon indicating the number of paintings associated with that painter. The figure’s basic information (e.g. name, birth year, death year, nickname, status) will be displayed when the art historian hovers over the figure icon, while all of their events will be shown in the table below. For the types of historical events that facilitate further analysis by art historians, we have coded them separately with different colors (Fig. 5). Art historians can click on the lines connecting the figures to view the events associated with them in the table below. The thickness of the lines represents the number of events, providing a visual indication of the strength of the relationship between the figures.
5.2.2 Cohort View
Cohort analysis is essential for art historians to understand the relationships and interactions between historical figures [68, 69]. Matrix provides a comprehensive and efficient way to visualize these complex relationships and identify patterns. The Cohort View (Fig. 3E2-2) displays the relationship between the sealer and inscriber of a painting. We also support art historians in adding new figures to the matrix for further exploration. The type and shade of color are used to encode the event category and quantity.
5.3 Similar Painting View
Art historians hope to discover potential painting apprenticeship relationships through the similarity of the painting styles (Fig. 3E3). The approach outlined in Section 4.1.2 was utilized to obtain data on paintings that exhibit similarities to the original painting. Meanwhile, some paintings are similar to each other in terms of theme, so we use Term Frequency-Inverse Document Frequency (TF-IDF) [60] model to obtain similar painting data. In order to facilitate art historians in comparing paintings, we set the view position to be close to the original painting and sort the paintings according to the author’s birth year. When an art historian is interested in a similar painting, they can click on the [Uncaptioned image]Handscroll Feature View button to start a new exploration of the painting. Similarly, if they are interested in the author of a similar painting, they can click on either the [Uncaptioned image]Linked Data View or [Uncaptioned image]Painting Biography View button to explore that painter.
5.4 Painting Biography View
The Painting Biography View demonstrates the order of collectors and appreciators of the painting along a timeline (Fig. 3E4) (R3).
Mapping figure type with shape and color. The figures displayed in the view are divided into three types (Fig. 5). The painting’s author is represented by a circular seal. The collectors and appreciators are represented by a square seal with the same color code as in Fig. 3E1-1. Figures complemented by art historians have a dashed border.
Visualizing the life span with handscroll. We use the handscroll as a metaphor for the figure’s life span (Fig. 5), which term the life-handscroll. Above the life-handscroll, we use black blocks to encode inscriptions with exact time, while the size of the blocks encodes the number of words in each inscription. Below the life-handscroll, we use red squares to encode seals with the exact time. The untimed inscriptions and seals are placed on the left side of the life-handscroll, and the number in the red square represents the number of seals. We encode common events as a line with a clickable diamond and a stacked bar chart is designed to facilitate the analysis of the distribution and types of events among figures, which can provide valuable insights into the social and cultural contexts of the figures’ lives.
Figure 5: Each figure icon in the Painting Biography View encodes the figure’s basic information and its comprehensive level, and related events between figures.
Displaying figure importance with hierarchy. Under the guidance of art historians, we have developed a comprehensive evaluation metric for determining the importance of figures. To evaluate a figure’s importance score $S$, we consider three key factors:
  • Painting relevance ($R$). $R=w_{0}r_{0}+w_{1}\log r_{1}+w_{2}\log r_{2}+w_{3}\log r_{3}.$ It contains four aspects, a) whether the figure is a pioneer or representative of a painting school, and the importance $r_{0}$ of the painting scroll, b) the number $r_{1}$ of his paintings, c) the number $r_{2}$ of his paintings being appreciated, and d) the number $r_{3}$ of paintings he appreciated. The coefficients $w_{0}$, $w_{1},w_{2}$, and $w_{3}$ are used to balance the four terms. We reduce the weight of Qing Dynasty data in our analysis for balance, due to its excess of paintings.
  • Degree of discussion ($D$). $D=\log d.$ The number of times $d$ the painter is mentioned in ancient literature [1].
  • Figure identity ($I$). $I=\sum_{i\in\mathcal{I}}s_{i}.$ where $\mathcal{I}\subseteq$painting collector, literati, official} is the identity set. The corresponding $s_{i}$ for painting collectors and literati are both assigned a value of 10, while the $s_{i}$ for officials ranges from 0 to 20, depending on their official positions.
The complete score of the figure is finally defined by $S=\lambda_{1}R+\lambda_{2}D+\lambda_{3}I,$ where coefficients $\lambda_{1},\lambda_{2}$, and $\lambda_{3}$ determine the weights of factors. All of the coefficients are set as 1 by default and can be adjusted by the experts according to their preference.
Interaction. We provide art historians an interactive way to integrate events into painting biographies based on their research interests (T5). Painting biographies are categorized into three trust levels: painting (analysis of the painting’s content), reasoning (based on the painting and related external information), and recommendation (automatic recommendation outside of the painting) (T6). Art historians can select the appropriate level button as per their requirements.
5.5 Uncertainty Data View
Because of the inherent characteristics of historical data, a huge number of historical figures lack comprehensive information regarding birth and death dates. To ensure the integrity of biographies, such uncertain data has been retained in the Uncertainty Data View (Fig. 3E5).
6 Evaluation
In this section, we show the usefulness and usability of ScrollTimes through two case studies and interviews with experts.
6.1 Case studies
We invited the art historians mentioned in Section 3 to freely explore ScrollTimes according to their research interests. We encouraged them to use the think-aloud protocol, and we recorded their usage of the system in two cases.
Figure 6: Exploration Process. The scene and the figures in the painting “The Elegant Gathering in the Western Garden” have been explored (A). A1 recognizes the main figures’ information through the word cloud (A4) and the inscriptions (B), and the relationship network between the figures indicate that they have formed a close cohort (C). A1 then explores all figures’ life-scroll to study this idealized gathering (D). Through the Similar Painting View, A1 finds that “the Elegant Gathering” becomes an influential theme in future generations (E).
6.1.1 Case 1: Autumn Colors on the Que and Hua Mountains
A4 has a particular fondness for the paintings of Zhao Mengfu, an outstanding painter in the Yuan Dynasty. A4 chooses “Autumn Colors on the Que and Hua Mountains”, one of the most extraordinary paintings by Zhao, to dig deep into the painting’s biography.
What does the painting depict? By rotating the circularly-projected handscroll, A4 put the core painting to the top-center of vision (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA). Consequently, there are mountains with distinct shapes (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA1). According to the hint through the word cloud below (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA2), it is revealed that the mountain on the left is Que Mountain, while the mountain on the right side represents Huabuzhu Mountain. It is exactly where the origin of the painting’s title comes from.
How was the painting created? To gain in-depth insights into the underlying motivation for its creation, A4 subsequently analyzed the inscription left by the painter Zhao Mengfu (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB). It recorded that Zhao Mengfu resigned from his official position in Qizhou (Jinan City, Shandong Province, China) in 1295 (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB1). In fact, the linked external dataset CBDB also contains a recorded event corresponding to the description, which, however, lacks the annotations of starting and ending years (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB2). After combining the two pieces of information, A4 proposed that the missing data could be filled in. It turns out that in-depth analysis based on linked data can even feed back into the complementary of the linked external databases. Zhao Mengfu’s inscription also showed that this painting was created for a figure named Gong Jin (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB3), whose frequent appearance in the word cloud also caught A4’s interest. However, A4 was not familiar with this figure. By leveraging the Linked Data View, A4 found that Gong Jin was an alias of Zhou Mi, a celebrated poet of the Southern Song Dynasty known by other aliases Bianyang Laoren (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB4). These terms also emerged frequently in the word cloud, which is now known as other references to Zhou Mi. Further inquiry into the historical context of the painting revealed that Zhou Mi and Zhao Mengfu were indeed close friends (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryB4).
Who collected and appreciated this painting? After creation, a famous painting will be widely collected and appreciated by future generations. A4 utilized the Biography View to trace the painting’s biography from ancient times to the present (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC).
The Biography View displays the involved historical figures who left seals and inscriptions on the handscroll. As A4 noticed, Zhou Mi was absent since he did not leave any seal or inscription on the painting. However, as previously mentioned, the painting was created for Zhou Mi, which means Zhou Mi still plays a vital role in the painting’s biography. A4 thus added Zhou Mi to the Biography View manually (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC1). The Biography View implies that this painting has been passed down through the Yuan, Ming, and Qing dynasties and has been collected or appreciated by fifteen historical figures. Accordingly, a few connections of common events exist among those figures, where the [Uncaptioned image]Political and [Uncaptioned image]Academic relations are most representative. It is worth noting that three Chinese emperors also left seals on the painting, with Emperor Qianlong leaving up to 33 seals and 9 inscriptions. A4 found something interesting: Emperor Qianlong pointed out a mistake in the painter’s inscription. Zhao Mengfu incorrectly stated the position between Que Mountain and Hua Mountain (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC2, B5). However, after examining the content of the inscriptions left by Emperor Qianlong in detail, A4 found that they all revolved around Qianlong’s experience of visiting Qizhou and viewing the Que and Hua Mountains (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC3). His repeated inscriptions on the same event did not provide much valuable information. Therefore, A4 chose to lower his importance (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryC4). From the above steps, A4 safely drew the collection and appreciation path of this painting, which serves as the most essential part of the painting biography. Nevertheless, A4 tried to investigate the painting genres to complement the biography.
Who is influenced by this painting? The painting style and skills will be passed on, affecting other people, which is also a crucial part of the painting’s biography. Specifically, A4 found that a certain similar painting style occurs in other paintings from Huang Gongwang (ScrollTimes: Tracing the Provenance of Paintings as a Window into HistoryA3), Ni Zan, Wang Hui, and Tang Yin, etc. It inspired A4 to take an in-depth look at the Linked Data View, and A4 confirmed that Huang Gongwang did have a teacher-student relationship with Zhao Mengfu. Therefore, A4 added Huang Gongwang to the biography. So far, A4 indicated that he had a sufficient and comprehensive understanding of this painting, and had constructed its corresponding biography. As A4 stated, this handscroll, which has lived through three dynasties over for 700 years, has been connected to at least 20 influential figures. This handscroll is like a time capsule that holds many records of history.
6.1.2 Case 2: Conversations span the millenniums
A1 is an art historian who studies figure paintings, and he has a particular fondness for paintings with the theme of “Elegant Gathering”. To learn more about this theme, A1 used ScrollTimes to perform a thorough analysis of the masterpiece painting “The Elegant Gathering in the Western Garden” to better comprehend the perspective it portrays and the story surrounding it.
A close cohort of literati. A1 rotated the painting circle and observed this painting, in which the figures waved their writing brushes and painted, recited poems and lyrics, and meditated and received Zen instruction (Fig. 6A1). A1 looked at the seals and inscriptions extracted and identified from the painting, which didn’t arouse his research interest. However, Qianlong left many inscriptions providing important information, that is, the names of the figures in the painting. A1 viewed the word cloud and the inscriptions to further explore the figures in the painting and the story behind their gathering in the Western Garden. Combined with the word cloud (Fig. 6A4), A1 recognized that the main figures in the painting include Wang Shen, Su Shi, Su Zhe, Huang Tingjian, etc. The inscriptions show detailed information about the figures in the painting, and A1 learned that the story is that Wang Shen invited a painter to make records of the incident where he and his friends gathered in the Western Garden (Fig. 6B). The figures in the painting include well-known Su Shi and Huang Tingjian, but A1 was unfamiliar with others, so he placed these figures in the Linked Data View to explore related information. The Linked Data View revealed that every figure in the painting had affairs with Su Shi. The relationship network between the figures indicated that they had formed a close cohort (Fig. 6C). Further, through event color classification, it can be seen that most of the events occurring within the cohort were [Uncaptioned image]Academic events. Among them, there were also [Uncaptioned image]Paint interactions between Li Gonglin and Li Zhiyi, Wang Shen and Huang Tingjian. A1 realized that this painting shows a close cohort of literati with Su Shi as the core. This cohort has strong academic resources, promoting literary and artistic innovation since the mid and late Northern Song Dynasty.
An imaginary gathering. Analysis of the close cohort aroused A1’s curiosity, so A1 raised a question: Is it possible for so many scholars and celebrities to gather together at the same time? To explore this answer, A1 added figures to the Biography View to examine the possibility of the intersection of this cohort of literati (Fig. 6D). The Biography View shows the birth and death dates of all the figures on the painting, all of whom lived in one era. A1 analyzed, “In terms of time, the figures in the painting were given the opportunity to gather on one day.” A1 indicated that this form of the gathering was reminiscent of the masterpiece of Western art history, Raphael’s “The School of Athens”. Raphael did not dwell on the real Plato Academy in history when he painted it; instead, he painted the ideal “the School of Athens” [56]. So A1 guessed, “This form of expression may coincide with this painting, although the time allows this gathering, but whether it is a real gathering needs further historical evidence to support.”
A legacy across the ages. In order to explore the impact on other paintings and the potential painting apprenticeship relationships between paintings, A1 used the Similar Painting View to observe similar paintings (Fig. 6E). Among theme similar paintings, A1 found three paintings with the same theme of “Elegant Gathering”, across the Song, Yuan, and Ming dynasties, painted by Li Gonglin, Liu Songnian and Tang Yin (Fig. 6E). A1 indicated that “Elegant Gathering” was a famous theme in the history of painting, so it was natural to find some paintings with similar themes. This showed that later generations did not simply learn and imitate this painting in terms of painting technique and style, but more importantly, they revered and admired the literati collective centered on Su Shi. People’s longing for the elegance, spirit, and style of the literati and scholars of the Song Dynasty had led to the repeated imitation and interpretation of “Elegant Gathering” by later generations of painters, and the creation is not only a manifestation of their cultural aesthetics but also the ideal world they constructed. Finally, A1 constructed a painting biography by exploring the stories of the figures and gained a comprehensive understanding of “The Elegant Gathering in the Western Garden.”
6.2 Art Historian Reviews
We evaluated the efficacy and usability of ScrollTimes by interviewing experts ( Section 3) online, who were not co-authors of this manuscript.
Procedure. Each 90-minute interview consisted of a 5-minute consent and demographic questionnaire, a 20-minute introduction to our motivation and system design, a 45-minute free exploration of ScrollTimes, and a 20-minute evaluation of its effectiveness and usability.
Effectiveness of the system. All experts agreed that ScrollTimes would enhance their research. A1-A5 found it particularly useful for improving efficiency in seal identification and inscription analysis. A4 mentioned their current workflow was still in the “manual workshop” stage, involving searching different platforms to confirm the seal owner, leading to a time-consuming and inefficient process. This tedious and repetitive work adversely affects their research efficiency and enthusiasm. The automatic method provided by ScrollTimes can alleviate the laborious seal extraction process, making it a productivity tool that allows them to focus on other analytical tasks. A3 found that the combination of multiple databases saves time searching for background information on figures in paintings, and ScrollTimes can help them quickly find research entry points. A4 and A5 suggested using ScrollTimes as an auxiliary tool in the classroom of art history, improving teaching efficiency and student learning interest. E1 and E2 found the social relationship information reflected in the inscriptions helpful. E1 suggested the inscriptions could be a reliable data source to supplement the CBDB database, verifying existing data and expanding the data source, breaking down barriers between painting and text data.
Usability of the system. All experts agreed that the view designs are reasonable, and the exploration process is clear. A3-5 found the circular design of the handscroll (Fig. 3E1) to be “eye-catching.” Although the circular design may compromise the original aesthetic beauty of the handscroll, A4 believed that it allows viewers to see the whole picture of the handscroll at a glance, and the interaction with the circular design does not lose any original information. A3 and A5 approved of the word cloud design in the circular design, believing it provides a good summary of seal and inscription information and can inspire analysis. The comparative presentation in the Seals and Inscriptions View (Fig. 3E1-3) enhances the results’ trustworthiness. The Linked Data View (Fig. 3E2) provides rich person-to-person association information that the experts welcomed. While the Cohort View (Fig. 3E2-2) was familiar to E1-2, art historians expressed the need for some learning costs. A1 and A4 found the Biography View to be the most impressive, stating that “using a familiar scroll as a metaphor for life is excellent.” The coding on the life-scroll is rich but not difficult to understand.
Suggestions. Art historians provided valuable suggestions based on their research interests. A4 suggested enhancing the comparison of painting biography by supporting the comparison of multiple painting biographies based on the same theme. A1-A3 suggested incorporating more ancient book data and displaying the sources of the data.
7 Discussion
This section summarizes the lessons learned from working with art historians (A1-A5) and CBDB editors (E1 and E2), as well as the limitations and future work of ScrollTimes.
7.1 Design Implications
Open-mindedness. In addition to being a carrier of historical information, the handscroll is also a traditional work of art. Out of respect for ancient artworks, art historians typically favor the original form of the paintings and tend to view new technological solutions for processing them as potentially damaging, such as the circle transformation of the handscroll, leading to doubts and concerns. However, once our handscroll processing method is completed and presented, art historians realize the immense benefits of the new approach, evincing a high degree of interest and appreciation, and exhibiting readiness to embrace and adopt this technological solution. When engaging in interdisciplinary research, it is crucial to recognize that our new proposals may encounter resistance due to violating empirical habits. Therefore, it is important to comprehensively instruct and present the potential advantages of the new methods to collaborative researchers so that they can gain a thorough comprehension, ultimately appreciating the proposed solution.
Data credibility. Art historians place great importance on the reliability of data sources, as unreliable or unclear data sources empirically lead to unreliable or even wrong conclusions. This necessitates a substantial amount of effort to verify the data, diverting attention away from more in-depth research pursuits. In contrast, the proposed ScrollTimes system has undergone comprehensive verification and supplementation from multiple sources. Moreover, the system also includes detailed information on data sources, origins, and source database links. Such features can significantly bolster art historians’ trust in the data, rendering them more amenable to utilizing the system.
Data complementarity. Most of the databases and systems operate unidirectionally, with data providers solely responsible for managing and distributing data to researchers. However, our system features a bidirectional approach that enables researchers not only to access data from the database but also to contribute additional information through analysis to enhance the database. This approach facilitates mutual data complementation and reinforcement, resulting in a more robust and comprehensive database.
Expansion of research fields. Our work establishes a workflow in the field of handscrolls that leverages multiple models to process and integrate diverse data sources to gain profound insights. This approach can be expanded to research other domains of cultural heritage. For example, in the investigation of ancient architectural heritage, in addition to the information from the architecture itself, various multimodal sources such as related books, paintings, poems, etc. can also be used for verification and supplementation.
7.2 Limitations and Future Work
Aesthetic exploration. The ScrollTimes system has primarily focused on exploring the historical value of handscrolls. Still, art historians have pointed out that we should also pay more attention to the aesthetic value of these works and provide analysis protocols on the painting style, brushwork, etc. Although we have included similar paintings in core painting and subject directions, art historians have expressed the need for more detailed similarity calculations based on other artistic styles, such as brushwork and ink, to enhance their analysis and research.
Data source expansion. Although our ScrollTimes system has already integrated numerous databases, we are eager to incorporate additional data sources, particularly those related to aesthetic documents, such as “Empty and Full: The Language of Chinese Painting” [15] and “The Chinese on the Art of Painting” [58] and so on. However, these data are merely recorded as raw texts, lacking a systematic and structured database format, which makes it challenging to use them for verification and supplementation.
Painting biography comparison. The ScrollTimes system does not provide an intuitive comparison of painting biographies, and art historians have suggested that comparing the biographies of different paintings can reveal cohort-level cultural heritage activities, thereby enabling analysis and comparisons of heritage styles. Those shortcomings highlight opportunities for improvement in future work.
8 Conclusion
To thoroughly study handscrolls, an art form that contains rich historical information, we present a visual analysis system called ScrollTimes, which tracks the origins and information contained in handscrolls by linking multiple data sources. We extract, verify, and supplement the elements and information contained in handscrolls, and create an interactive biography to further reveal their essence. Our approach offers a window into history, fostering a holistic understanding of handscroll provenance and historical significance. Two case studies and expert interviews demonstrate the usefulness and effectiveness of our system. We have summarized the lessons learned from developing ScrollTimes and believe that these implications will guide the system that is more comprehensive in functionality, richer in data sources, and more effective as a tool for art historians.